What is the slope of the line with the equation 2x+3y+6=0?
Answer by lwsshak3(11628)

What is the slope of the line with the equation 2x+3y+6=0?
**
Standard form of equation for a straight linee: y=mx+b, m=slope and b=y-intercept.
2x+3y+6=0
Rewrite in standard form:
3y=-2x-6
y=-2x/3-3
slope m=-2/3